The average length of time for student to register summer classes at a certain university has been 50 minutes with a standard deviation of 10 minutes. A new registration procedure using modern computing machine is being tried. If a random sample of 12 students had a n average registration of 42 minutes with a standard deviation of 11.9 minutes under the new system, test the hypothesis that the population mean is now less than 50 minutes. (critical value is +/- 1.796)


1
Expert's answer
2022-05-24T17:45:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge50"

"H_1:\mu<50"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=11" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.796."

The rejection region for this left-tailed test is "R = \{t:t<-1.796\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{42-50}{11.9/\sqrt{12}}\approx-2.3288"

Since it is observed that "t=-2.3288<-1.796=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=11" degrees of freedom, "t=-2.3288," is "p= 0.019977\n," and since "p= 0.019977
<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 50, at the "\alpha = 0.05" significance level.